OpenStudy (anonymous):

8^4/3 express the number in the form a/b where a and b are integers

OpenStudy (anonymous):

Which one is it? \[8^4\div3\] or \[8^(4\div3)\]

OpenStudy (anonymous):

8 to the 4/3 power

OpenStudy (anonymous):

or the power of 4/3 however you say it

OpenStudy (anonymous):

Ok. then 8^(4/3) = (2^3)^(4/3) = 2^4 = 16 or 16/1

OpenStudy (anonymous):

\[8^(4\div3) = (2^3)^(4\div3) = 2^4 = 16\]

OpenStudy (anonymous):

that's what i got and it's wrong

OpenStudy (anonymous):

so 16/1 is wrong?

OpenStudy (anonymous):

I mean, did you get 16/1 as wrong? or 16 as wrong?

OpenStudy (anonymous):

nevermind i didn't put it over one

OpenStudy (anonymous):

Yo! There you go.